OpenStudy (anonymous):

8 times the square of a nonzero number is equal to 96 times the number. What is the number?

OpenStudy (anonymous):

please help /.\

OpenStudy (anonymous):

a. 24 b.12 c.96 d.4

OpenStudy (anonymous):

\( 8x^2 = 96x \) <-------- That is the equation. Can you solve?

OpenStudy (anonymous):

8x^2 = 96x 8x^2 - 96x = 0 (use the quadratic equation to solve for x)

OpenStudy (anonymous):

/.\ haha @MDoodler i suck at math sooo /.\ i cant really

OpenStudy (anonymous):

Can you solve \( 8x^2 -96x = 0 \) ?

OpenStudy (anonymous):

12 24 96 4 those are the choices they give me

OpenStudy (anonymous):

Can you solve the last equation I gave you or attempt to solve it and tell me your answer?

OpenStudy (anonymous):

/.\ i tryyed i got 24

OpenStudy (anonymous):

The answer is 12

OpenStudy (anonymous):

You want to know how I got the answer?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

\[ \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a}\]

OpenStudy (anonymous):

I used this http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

use the quadratic equation I just posted

OpenStudy (anonymous):

The link I gave you shows you how to use the quadratic equation

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

Also, to check if you did it right, you can use this link http://www.wolframalpha.com/input/?i=8x%5E2+-96x+%3D+0 It will solve your work for you just by entering the equation.